OpenStudy (anonymous):

Determine the coordinates of the absolute maximum and minimum values of the function below for the given interval. f(x) = 3xe^(-x) For interval -1 less than or equal to x less than or equal to 3

OpenStudy (idku):

\[[-1,3]\]

OpenStudy (idku):

\[f'(x)=?\]

OpenStudy (anonymous):

3e^(-x)+3x*-e^(-x)

OpenStudy (idku):

yes, or \[3e^{-x}-3xe^{-x}=3e^{-x}~(1-x)\]

OpenStudy (anonymous):

Thank you I think this is where I was messing up!

OpenStudy (idku):

now, critical numbers are those that: 1. where f'(x) is undefined, iFF part of f(x) domain. 2. x-values when f'(x)=0 3. closed interval boundaries.

OpenStudy (idku):

There are no restrictions for x in f(x) or f'(x). set f'(x)=0

OpenStudy (idku):

and solve for x.

OpenStudy (idku):

\[0=3e^{-x}(1-x)\]

OpenStudy (idku):

x=?

OpenStudy (anonymous):

1

OpenStudy (idku):

yes

OpenStudy (anonymous):

I was messing up getting the critical point out of the derivative thanks a lot :-)

OpenStudy (idku):

so, your interval boundaries (closed - i.e included in your part of the function) are -1 and 3 so critical numbers are {-1, 1, 3}

OpenStudy (idku):

now, find f(-1) f(1) and f(3) and tell me which one is the minimum and the maximum

OpenStudy (anonymous):

min at f(-1) and max at f(1)

OpenStudy (idku):

check, are you sure ?

OpenStudy (anonymous):

Yeah!

OpenStudy (idku):

f(x) = 3xe^(-x) f(3)=3(3)e^(-3)=9/e^3=0.44 f(1)=3(1)e^(-1)=3/e=1.1 f(-1)=3(-1)e^(--1)=-3e^2 yes, I was looking at the derivative... you are right. sorry

OpenStudy (idku):

so, min is (-1,f(-1)) max is (1,f(1))

OpenStudy (idku):

bye!

OpenStudy (anonymous):

Thank you a ton!